Use the substitution method to solve the linear system given. Express the answer as an ordered pair of the form (x,y)
x + y = 5
2x + y = 6
I need these answers to be as straightforward as possible. No shown work is needed. *Thank you for all simply put answers*

Found 2 solutions by texttutoring, richard1234:
Answer by texttutoring(324)

x+y=5
2x+y=6

Let`s isolate for x in equation 1:

x = 5-y

Now let`s substitute this value into 2x+y=6:

2(5-y) + y = 6
10 - 2y + y = 6
10 - y = 6
10 - 6 = y
y = 4

Now find x:

x=5-y
x=5-4
x=1

So the answer, as an ordered pair is: (1,4)

Answer by richard1234(7193)

I prefer you show your work to avoid losing points on a test and so that your answer/proof is backed up. I'll show the first step, then you should work from there.

Note that 2x + y = (x + y) + y = 6. From this you can easily solve for y, then plug it in to solve for x.